OpenStudy (anonymous):

ln(x+3) Find the domain.

OpenStudy (anonymous):

I thought it was (-3,inf) because the domain must be greater than zero, right?

OpenStudy (anonymous):

Yes \[(-3nf)\]

OpenStudy (anonymous):

but this is wrong

OpenStudy (anonymous):

If it isn't (-3,inf) what is it and why?

OpenStudy (anonymous):

it is

jimthompson5910 (jim_thompson5910):

The answer is (-3, inf) but maybe they want you to use something else for "inf" (like the actual infinity symbol) or maybe they want the answer in set-builder notation

OpenStudy (anonymous):

Well I am using WebAssign and I've used the infinity symbol. It still isn't working? :(

OpenStudy (anonymous):

on line course i bet, so there is probably some syntax error

OpenStudy (anonymous):

can you post the exact question? i have some familiarity with WebAssign

jimthompson5910 (jim_thompson5910):

oh it looks like ln(x) + 3

jimthompson5910 (jim_thompson5910):

so that would make the domain to be (0, inf)

OpenStudy (anonymous):

aaah i see the problem

OpenStudy (anonymous):

oops! That was my error, sorry

OpenStudy (anonymous):

\(\ln(x)+3\) is what it means not \(\ln(x+3)\)

jimthompson5910 (jim_thompson5910):

that's fine, I've seen plenty of cases where the system will write out ln x + 3 and actually mean ln(x+3) which is very confusing and frustrating

OpenStudy (anonymous):

While we are at it, how could I find the X-intercept on this graph? Clearly it doesn't start from the origin

jimthompson5910 (jim_thompson5910):

y = ln(x) + 3 0 = ln(x) + 3 ... plug in y = 0 solve for x

OpenStudy (anonymous):

because webassing is hypocritical it makes you write "sin(x)" but will itself write lnx

OpenStudy (anonymous):

lol, thanks guys! :)